Arrange the following solutions in (a) order of decreasing freezing point and in (b)
order of increasing boiling point: 0.10
1
Expert's answer
2020-02-12T05:48:54-0500

Well the question seems to be incomplete but let's see how to deal these type of questions

Suppose we are given

Q.)Arrange the following aqueous solutions in order of decreasing freezing points (lowest to highest temperature)"0.10 m\ Na_3PO_4,\ 0.35 m\ NaCl,\ 0.20 m\ MgCl_2,\ 0.15 m\ C_6H_{12}O_6 \ and \ 0.15 m\ CH_3COOH."


Solution

"\Delta T=i\times K_f\times m"

where i is the van't Hoff factor which is 4 for Na3PO4, 2 for NaCl, 3 for MgCl2, 1 for C6H12O6 and 1 for CH3COOH

basically here we have to see Larger the concentration of the concentration of the particles, smaller the freezing point.


0.10m Na3PO4

Na3PO4 → 3Na+ + PO43-

⇒ van't Hoff factor is 3 + 1 = 4

0.10 m * 4 = 0.40


0.35m NaCl

NaCl → Na+ + Cl-

⇒ van't Hoff factor = 1+1 = 2

0.35 * 2 = 0.70


0.20m MgCl2

MgCl2 → Mg2+ + 2Cl-

 ⇒ Van't Hoff factor = 1+2 = 3

0.20 * 3 = 0.60


0.15m C6H12O6

⇒ for non-ionic compounds in solution, like glucose (C6H12O6) , the van't Hoff factor is 1. They do not dissociate in water.


0.15 * 1 = 0.15


0.15m CH3COOH.

CH3COOH ⇄ CH3COO- + H+

 ⇒ Van't hoff factor ≈ 1<x<2

0.15 * 2 = 0.30 to 0.15


Larger the concentration of the concentration of the particles, smaller the freezing point.

0.15 m C6H12O6 > 0.15 m CH3COOH > 0.10 m Na3PO4 > 0.20 m MgCl2 > 0.35 m NaCl